# Online Optimization for Learning to Communicate over Time-Correlated Channels

### v1

@@ I Introduction
The widespread adoption of machine learning techniques in developing communication systems based on real-world data has sparked broad interest in recent years [1]. Machine learning algorithms have shown to be effective tools for various tasks such as channel estimation [2], equalization [3], coding [4], decoding [5], and other physical layer applications [1, 6]. Learning-based communication systems have showed impressive performance in their capacity to generalize effectively to unknown channels [1, 7].
@@
Figure 1: The procedure of learning decoder and codebook using online optimization methods.
@@
However, meeting the assumption of I.I.D. channels in reality is generally challenging, given that practical communication scenarios often involve time-varying channels with dynamic statistical properties. Among time-varying channels, a prominent example is the time-correlated channel [14, 15]. In real-world communication processes, there exist numerous instances of time-correlated channels. For example, user mobility usually leads to time-correlated fading channels in mobile communication [16, 17]. As a consequence of moderate user mobility, channel fading gains become interdependent between consecutive time slots, resulting in the well-known Markov fading channels [18, 19]. Hence, establishing generalization bounds for learning-based communication systems over time-correlated channels is challenging, as statistical learning theory, which significantly depends on the I.I.D. channel assumption, is not applicable in this scenario.
@@
In this paper, we will establish online optimization algorithms [20, 21] for learning communication systems over time-correlated channels with solid theoretical guarantees. The outlined online optimization procedure is depicted in Fig. 1. To be specific, we focus on two types of channels: 1) time-correlated fading channel with additive white Gaussian noise, where the fading distribution is unknown and time-correlated; and 2) time-correlated additive noise channel with an unknown and time-correlated noise distribution. For the time-correlated fading channel, the transceiver is equipped with a fixed codebook (constellation) and is tasked with optimizing the channel decoder using collected channel input-output pairs data. Besides, based on the proposed convex surrogate loss, we can regard such online channel decoder learning problem as an online convex optimization problem. In the case of time-correlated additive noise channels, the goal of transceiver becomes to select optimal codebooks during the whole communication process. We model this online codebook learning problem as a multi-armed bandit problem: the transceiver selects optimal codebooks from a predefined super-codebook using empirical data collected in real-time.
@@
  • For the first time, we formulate the task of learning channel decoder or codebook for communication over time-correlated channels as an online optimization problem without relying on the I.I.D. channel assumption, which is more in line with real communication scenarios.
@@
The remainder of the paper is structured as follows: In Section II, we present a review of prior related works on learning-based communication systems and online optimization theory. Section III introduces two distinct time-correlated channels and outlines the online optimization problem addressed in this study. Section IV and V detail the development of online optimization algorithms for learning channel decoder and codebook under these two time-correlated channels respectively. Section VI encompasses numerical simulations, while Section VII concludes the paper.
@@ II Related works
Online optimization can effectively model numerous online machine learning problems, encompassing online convex optimization [20, 27], prediction with expert advice [28], multi-armed bandit [29, 30] and more. This framework deals with a sequential decision problem, where a learner repeatedly takes actions within a feasible set and encounters a potentially adversarial loss function from the environment [20, 21]. During this decision-making process, the learner endeavors to devise an algorithm that minimizes regret, defined as the disparity between the total loss incurred by the learner and that of the best decision in hindsight. Departing from statistical learning theory, online optimization theory can furnish theoretical guarantees for algorithms without relying on the I.I.D. assumption. Recently, it has been noted that practical environments often exhibit predictable patterns, which the learner can leverage to achieve reduced regret [22, 24]. To pursue this objective, an optimistic Online Mirror Descent (OMD) framework has been proposed to harness environmental correlations for improving the performance of learned decisions [22, 24]. For instance, [23] employs optimistic OMD to solve the multi-armed bandit problem in predictable environments, while [31] utilizes it to address the stochastically extended adversarial online convex optimization problem.
@@ III-B Time-varying Channel Models
In this paper, we assume that the transmitter sends messages to the receiver over multiple rounds, with the total number of rounds denoted as $T$. We consider the problem of communication over time-varying channels defined below:
@@
where $X_{t}\in\mathbb{R}^{d}$ is a codeword which is chosen from a codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ with a uniform probability and $Y_{t}\in\mathbb{R}^{d}$ is the corresponding channel output in $t$-th round. $f_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a channel transformation, and $Z_{t}\in\mathbb{R}^{d}$ is a channel noise statistically independent of the input $X$ and the transformation $f_{t}$. In this paper, we assume that the codebook $C_{t}\in\mathcal{C}\subset(R^{d})^{M}$ and $\mathcal{C}$ is with the power constraint, i.e., $\mathcal{C}=\{C:\|\mathbf{x}^{j}\|_{2}\leq\gamma_{X},\forall j\in[M]\}$.
The statistical properties of channel transformation $f_{t}$ and channel noise $Z_{t}$ across different rounds are assumed to be different in this work. In other words, we consider time-varying channels and thus the distributions of $f_{t}$ and $Z_{t}$ can vary at each round during the communication process.
In this paper, we delve into two specific cases of the time-varying channel model described earlier, both of which hold high relevance in realistic communication scenarios. The first case involves a time-correlated fading channel, assuming that $f_{t}$ is a time-correlated linear transformation while the channel noise $Z_{t}$ is an independently and identically distributed (I.I.D.) Gaussian noise. The second case pertains a time-correlated additive noise channel, where $f_{t}$ is an identity transformation and $Z_{t}$ is a time-correlated additive channel noise.
@@ III-B1 Time-correlated Fading Channels
In the first considered channel model, we fix the codebook $C_{t}=C=\{\mathbf{x}^{j}\}_{j\in[M]},\forall t\in[T]$ for the transceiver at each round. The channel transformation is assumed to be the linear transformation and it is thus similar to the common fading channel in wireless communications [32, 33]. Hence, Eq. (1) becomes
@@
For this channel, we assume that the decoding rule is chosen from the class of nearest neighbor decoder with a linear kernel operating on the channel output, i.e., given a channel output $\mathbf{y}_{t}^{j}$, the index of the decoded codeword is selected as:
$j_{t}(\mathbf{y}^{j}_{t})\in\mathop{\arg\min}\limits_{j^{\prime}\in[M]}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}^{j}_{t}\|_{2},$ (3)
where $\mathbf{y}_{t}^{j}=\mathbf{H}_{t}\mathbf{x}^{j}+W_{t}$ and $\mathbf{G}_{t}\in\mathbb{R}^{d\times d}$ is a linear kernel operating on the channel output $\{\mathbf{y}^{j}_{t}\}_{j\in[M]}$.
Now we explain this choice of decoding rule. Initially, the selection of the nearest neighbor decoder is based on its optimality under the condition that $\mathbf{H}_{t}=\mathbf{G}_{t}$ and $W$ is Gaussian [32], but we do not assume that the channel satisfies this condition here [34]. Furthermore, the use of the linear kernel $\mathbf{G}_{t}$ on $\mathbf{y}_{t}$ shares similarities with the channel equalizer widely employed in communication systems [3, 32], aiming to approximate the channel gain $\mathbf{H}_{t}$ and minimize its influence on decoding. For simplicity, we refer to the linear kernel $\mathbf{G}_{t}$ as the channel decoder in the subsequent discussions.
We then define the $t$-th round expected error probability of the channel decoder $\mathbf{G}_{t}$ learned in the $t$-th round given that $M$ distinct codewords in $C$ are transmitted:
@@
where $j_{t}:=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}$.
Then, we consider conducting an online optimization procedure to identify the optimal channel decoder $\mathbf{G}_{t}$. The online optimization goal related to this motivation is to leverage channel output samples $\{\mathbf{y}_{t}^{j}\}_{j\in[M]},\forall t\in[T]$ generated at each round to construct the channel decoder $\mathbf{G}_{t}\in\mathcal{G}$ with minimal expected error probability for each round. In this paper, we assume that $\mathcal{G}=\{\mathbf{G}\in\mathbb{R}^{d\times d}:\|\mathbf{G}\|_{F}\leq D\}$.
Given that the transceiver is provided with channel output samples related to the channel gain $\mathbf{H}_{t}$ and the channel noise $W_{t}$ at each round, the empirical symbol error rate of the channel decoder $\mathbf{G}_{t}$ in the $t$-th round can be defined as
@@
Hence, the transceiver can utilize this symbol error rate to carry out the online optimization protocol for learning the channel decoder $\mathbf{G}_{t}$. We design the online convex optimization algorithms to learn $\mathbf{G}_{t}$ in Section IV.
@@ III-B2 Time-correlated Additive Noise Channels
where $Z_{t}\in\mathbb{R}^{d}$ is a time-correlated channel noise with the distribution $F_{Z_{t}}$, which is statistically independent of the input. Similarly, we do not make any assumptions on the distribution of $Z_{t}$, and $\{F_{Z_{t}}\}_{t\in[T]}$ is completely unknown to the transceiver. We suppose that the channel noise is time-correlated, i.e., $Z_{t}$ is dependent on $\{Z_{\tau}\}_{\tau\in[t-1]}$ from the previous round. We can also consider a first-order Markov noisy channel: $Z_{t+1}=Z_{t}+\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{d}$ is a random vector. This channel is pertinent to practical communication scenarios, where the presence of thermal noise in the receiver leads to gradual variations.
For this channel, we fix the channel decoder $\mathbf{G}_{t}=I$ for each round since the channel transformation $f_{t}$ is an identity function in this scenario. Similar to the time-correlated fading channel, we also choose the nearest neighbor decoding rule based on the Euclidean distance, i.e., given a channel output $\mathbf{y}_{t}^{j}$, the index of the decoded codeword is chosen as:
$j_{t}(\mathbf{y}^{j}_{t})\in\mathop{\arg\min}\limits_{j^{\prime}\in[M]}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}^{j}_{t}\|_{2}$ (7)
where $\mathbf{y}_{t}^{j}=\mathbf{x}_{t}^{j}+Z_{t},\forall j\in[M]$ and the codeword $\mathbf{x}_{t}^{j}$ is from the codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$. Notice that we do not assume that the time-correlated noise is Gaussian and thus this decoding rule is only selected for its simplicity [9, 34].
Analogously, we can define the expected error probability over the $t$-th round channel noise $Z_{t}\sim F_{Z_{t}}$ given that $M$ distinct codewords in $C_{t}$ are transmitted.
@@
where $j_{t}:=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}_{t}^{j}\|_{2}^{2}$. $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ denotes the codebook chosen in the $t$-th round.
The online optimization goal for this scenario is to utilize the channel output samples $\{\mathbf{y}_{t}^{j}\}_{j\in[M]},\forall t\in[T]$ for selecting the codebook $C_{t}\in\mathcal{C}\subset(R^{d})^{M}$ with minimal expected error probability at each round.
Based on channel output samples related to $Z_{t}$, the symbol error rate of codebook $C_{t}$ can be defined as
$\ell_{t}(C_{t}):=\frac{1}{M}\sum_{j=1}^{M}\mathbbm{1}\Big{\{}\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}^{2}<\|\mathbf{x}^{j}_{t}-\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}},$ (9)
Similarly, the transceiver can make use of this symbol error rate as the loss function to perform the online optimization procedure for learning the codebook $C_{t}$. We devise the multi-armed bandit algorithm to learn $C_{t}$ in Section V.

### v3

@@ I Introduction
The widespread adoption of machine learning techniques in developing communication systems based on real-world data has sparked broad interest in recent years [1]. Machine learning algorithms have shown to be effective tools for various tasks such as channel estimation [2], coding [3], decoding [4], and other physical layer applications [5]. Learning-based communication systems have showed impressive performance in their capacity to generalize effectively to unknown channels [6].
@@
However, meeting the assumption of I.I.D. channels in reality is generally challenging, given that practical communication scenarios often involve time-varying channels with dynamic statistical properties. Among time-varying channels, a prominent example is the time-correlated channel [13]. In real-world communication processes, there exist numerous instances of time-correlated channels. For example, user mobility usually leads to time-correlated fading channels in mobile communication [14, 15]. As a consequence of moderate user mobility, channel fading gains become interdependent between consecutive time slots, resulting in the well-known Markov fading channels [16, 17]. Hence, establishing generalization bounds for learning-based communication systems over time-correlated channels is challenging, as statistical learning theory, which depends on the I.I.D. channel assumption, is not applicable in this scenario.
(a) learn decoder via training sequence
(b) learn codebook via training sequence
Figure 1: The procedure of learning decoder and codebook via online optimization.
Figure 2: Transmission process of uplink communication from mobile user to base station.
In this paper, we will establish online optimization algorithms [18, 19] for learning communication systems over time-correlated channels with solid theoretical guarantees. Specifically, we consider optimizing the uplink communication process between a mobile user equipment acting as the transmitter and a base station serving as a receiver. One potential application of proposed algorithms is to enhance the performance of distributed learning conducted by multiple mobile devices within wireless networks [20, 21]. In each communication round, we consider that the transmitter sends a package to the receiver. This package is composed of two parts: the first part is a deterministic training sequence known to both the transmitter and receiver, and the second part consists of data symbols that carry the actual information intended for transmission. These data symbols are random and unknown to the receiver. We highlight that this training sequence serves a similar role of pilot signals used in channel estimation in mobile communication, which is essential for calculating the loss required for conducting online optimization. The outlined online optimization procedure is depicted in Fig. 1. Following the assumption widely adopted in the study on using online learning to optimize communication systems [22, 23], we presume that the channel remains constant during the transmission of both the training sequence and the data symbols within a single communication round. While the channel statistics varies across different rounds. Consequently, channel decoders and codebooks learned by training sequences can be utilized to improve the transmission performance of data symbols within the same round. The transmission process of uplink communication is shown in Figure 2.
@@
For the time-correlated channels studied in this paper, to be specific, we focus on two types of channels: 1) time-correlated fading channel with additive white Gaussian noise, where the fading distribution is unknown and time-correlated; and 2) time-correlated additive noise channel with an unknown and time-correlated noise distribution. For the time-correlated fading channel, the transceiver is equipped with a fixed codebook (constellation) and is tasked with optimizing the channel decoder using collected channel input-output pairs data. Besides, based on the proposed convex surrogate loss, we can regard such online channel decoder learning problem as an online convex optimization problem. In the case of time-correlated additive noise channels, the goal of transceiver becomes to select optimal codebooks during the whole communication process. Similar to [8], the codebook considered in this paper essentially refers to a modulation scheme, i.e., a constellation. We model this online codebook learning problem as a multi-armed bandit problem: the transceiver selects optimal codebooks from a predefined super-codebook using empirical data collected in real-time.
@@
  • For the first time, we formulate the task of learning channel decoder or codebook for communication over time-correlated channels as an online optimization problem without relying on the I.I.D. channel assumption, which is more in line with real scenarios.
@@
The remainder of the paper is structured as follows: In Section II, we present a review of prior works on learning-based communication systems and online optimization theory. Section III introduces two distinct time-correlated channels and outlines the online optimization problem addressed in this study. Section IV and V detail the development of online optimization algorithms for learning channel decoder and codebook under these two time-correlated channels respectively. Section VI encompasses numerical simulations, while Section VII concludes the paper.
@@ II Related works
Online optimization can effectively model numerous online machine learning problems, encompassing online convex optimization [18], prediction with expert advice [29], multi-armed bandit [30, 31] and more. This framework deals with a sequential decision problem, where a learner repeatedly takes actions within a feasible set and encounters a potentially adversarial loss function from the environment [18, 19]. During this decision-making process, the learner endeavors to devise an algorithm that minimizes regret, defined as the disparity between the total loss incurred by the learner and that of the best decision in hindsight. Departing from statistical learning theory, online optimization theory can furnish theoretical guarantees for algorithms without relying on the I.I.D. assumption. Recently, it has been noted that practical environments often exhibit predictable patterns, which the learner can leverage to achieve reduced regret [24, 26]. To pursue this objective, an optimistic Online Mirror Descent (OMD) framework has been proposed to harness environmental correlations for improving the performance of learned decisions [24, 26].
@@ III-B Time-varying Channel Models
In this paper, we assume that the transmitter sends packages to the receiver over multiple rounds, with the total number of rounds denoted as $T$. We consider the problem of communication over time-varying channels defined below:
@@
where $X_{t}\in\mathbb{R}^{d}$ denoting the random data symbol, is a codeword chosen from a codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ with a uniform probability. $Y_{t}\in\mathbb{R}^{d}$ is the corresponding channel output in $t$-th round. $f_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a channel transformation, and $Z_{t}\in\mathbb{R}^{d}$ is a channel noise statistically independent of the input $X$ and the transformation $f_{t}$. In this paper, we assume that the codebook $C_{t}\in\mathcal{C}\subset(\mathbb{R}^{d})^{M}$ and $\mathcal{C}$ is with the power constraint, i.e., $\mathcal{C}=\{C:\|\mathbf{x}^{j}\|_{2}\leq\gamma_{X},\forall j\in[M]\}$.
The statistical properties of channel transformation $f_{t}$ and noise $Z_{t}$ across different rounds are assumed to be different in this work. In other words, we consider time-varying channels and thus the distributions of $f_{t}$ and $Z_{t}$ can vary at each round during the communication process.
In this paper, we focus on two specific cases of this time-varying channel model, both of which hold high relevance in realistic scenarios. The first case involves a time-correlated fading channel, assuming that $f_{t}$ is a time-correlated linear transformation while the channel noise $Z_{t}$ is an IID Gaussian noise. The second case pertains to a time-correlated additive noise channel, where $f_{t}$ is an identity transformation and $Z_{t}$ is a time-correlated additive channel noise.
@@ III-B1 Time-correlated Fading Channels
In the first considered channel model, we fix the codebook $C_{t}=C=\{\mathbf{x}^{j}\}_{j\in[M]},\forall t\in[T]$ for the transceiver at each round. The channel transformation is assumed to be the linear transformation and it is thus similar to the common fading channel in wireless communications [32]. Hence, Eq. (1) becomes
@@
For this channel, we assume that the decoding rule is chosen from the class of nearest neighbor decoder with a linear kernel operating on the channel output, i.e., given a channel output $\mathbf{y}_{t}$, the index of the decoded codeword is selected as:
$j(\mathbf{y}_{t})\in\mathop{\arg\min}\limits_{j^{\prime}\in[M]}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}\|_{2},$ (3)
where $\mathbf{y}_{t}=\mathbf{H}_{t}X_{t}+W_{t}$ and $\mathbf{G}_{t}\in\mathbb{R}^{d\times d}$ is a linear kernel operating on the channel output.
Now we explain this choice of decoding rule. Initially, the selection of the nearest neighbor decoder is based on its optimality under the condition that $\mathbf{G}_{t}=\mathbf{H}^{-1}_{t}$ and $W_{t}$ is Gaussian [32], but we do not assume that the channel satisfies this condition here [33]. Furthermore, the use of the linear kernel $\mathbf{G}_{t}$ on $\mathbf{y}_{t}$ shares similarities with the channel equalizer widely employed in communication systems [32], aiming to approximate the channel gain $\mathbf{H}_{t}$ and minimize its influence on decoding. For simplicity, we refer to the linear kernel $\mathbf{G}_{t}$ as the channel decoder in the subsequent discussions.
As shown in Fig 1 (a), in each round, the transmitter sends both a deterministic training sequence including $M$ distinct codewords from $C$ and random data symbols to the receiver. We first define the $t$-th round expected error probability of learned $\mathbf{G}_{t}$ for $M$ codewords in the training sequence as:
@@
where $j_{t}:=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|_{2}^{2}$ and $\mathbf{y}_{t}^{j}=\mathbf{H}_{t}\mathbf{x}_{t}^{j}+W_{t},\forall j\in[M]$.
In this work, following the coherence time assumption considered in [23], we also consider that the training sequence and data symbols experience the same channel gain during a communication round. Consequently, the decoder learned by the training sequence can enhance the transmission performance of data symbols. This assumption is reasonable, as it aligns with the common practice of using similar assumptions for pilot signals in mobile communication [22, 34]. Besides, considering that data symbols are uniformly sampled from the codebook, we notice that $\mathbb{E}_{X_{t}}[Pr\{\hat{X}_{t}\neq X_{t}|X_{t};\mathbf{G}_{t}\}]=\mathbb{P}_{t}(\mathbf{G}_{t})$, where $\hat{X}_{t}$ represents the prediction of the channel output transmitted by data symbol $X_{t}$. Based on it, the empirical performance of $\mathbf{G}_{t}$ on the training sequence can be utilized to execute online optimization for identifying the optimal channel decoder on data symbols. The online optimization goal related to this motivation is to leverage channel output samples $\{\mathbf{y}_{t}^{j}\}_{j\in[M]},\forall t\in[T]$ transmitted by $M$ codewords in the training sequence at each round to construct the channel decoder $\mathbf{G}_{t}\in\mathcal{G}$ with minimal expected error probability for each round. In this paper, we assume that $\mathcal{G}=\{\mathbf{G}\in\mathbb{R}^{d\times d}:\|\mathbf{G}\|_{F}\leq D\}$.
Given that the transceiver is equipped with channel output samples transmitted by $M$ codewords in the training sequence related to the channel gain $\mathbf{H}_{t}$ and the channel noise $W_{t}$ at each round, we can calculate the empirical symbol error rate of the channel decoder $\mathbf{G}_{t}$ in the $t$-th round using the training sequence, which includes $M$ distinct codewords as:
@@
Based on the above analysis, we notice that if a learned decoder $\mathbf{G}_{t}$ shows lower $\mathbb{P}(\mathbf{G}_{t})$ for the training sequence, it can improve the transmission performance of data symbols. Hence, the transceiver can utilize the symbol error rate defined in Eq. (5) to carry out the online optimization protocol for learning the channel decoder $\mathbf{G}_{t}$ to enhance the transmission performance of data symbols. We design the online convex optimization algorithms to learn $\mathbf{G}_{t}$ in Section IV.
@@ III-B2 Time-correlated Additive Noise Channels
where $Z_{t}\in\mathbb{R}^{d}$ is a time-correlated channel noise with the distribution $F_{Z_{t}}$, which is statistically independent of the input. Similarly, we do not make any assumptions on the distribution of $Z_{t}$, and $\{F_{Z_{t}}\}_{t\in[T]}$ is completely unknown to the transceiver. We suppose that the channel noise is time-correlated, i.e., $Z_{t}$ is dependent on $\{Z_{\tau}\}_{\tau\in[t-1]}$ from the previous round. We can also consider a first-order Markov noisy channel: $Z_{t+1}=\sqrt{1-\mu_{t}}Z_{t}+\sqrt{\mu_{t}}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{d}$ is a random vector and $\mu_{t}\in[0,1]$. This channel is pertinent to practical communication scenarios, where the presence of thermal noise in the receiver leads to gradual variations.
For this channel, we fix the channel decoder $\mathbf{G}_{t}=I$ for each round since the channel transformation $f_{t}$ is an identity function in this scenario. Similar to the time-correlated fading channel, we also choose the nearest neighbor decoding rule based on the Euclidean distance, i.e., given a channel output $\mathbf{y}_{t}$, the index of the decoded codeword is chosen as:
$j(\mathbf{y}_{t})\in\mathop{\arg\min}\limits_{j^{\prime}\in[M]}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}_{t}\|_{2},$ (7)
where $\mathbf{y}_{t}=X_{t}+Z_{t}$ and the data symbol $X_{t}$ is uniformly sampled from the codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$. Notice that we do not assume that the time-correlated noise is Gaussian and this decoding rule is selected for its simplicity [8, 33].
Analogously, as shown in Fig 1 (b), the transmitter sends $M$ distinct codewords in the deterministic training sequence and random data symbols to the receiver in each round. We also define the expected error probability over the $t$-th round channel noise $Z_{t}\sim F_{Z_{t}}$ for $M$ distinct codewords in the training sequence as
@@
where $j_{t}:=\arg\min_{j^{\prime}\in[M]\setminus j}\|\mathbf{x}_{t}^{j^{\prime}}-\mathbf{y}_{t}^{j}\|_{2}^{2}$ and $\mathbf{y}_{t}^{j}=\mathbf{x}_{t}^{j}+Z_{t},\forall j\in[M]$. $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ denotes the codebook chosen in the $t$-th round.
Similarly, we presume that the training sequence and data symbols share the same noise statistics during a communication round in this scenario. The codebook learned by the training sequence can thus contribute to improving the transmission reliability of data symbols. Since data symbols are uniformly sampled from the learned codebook $C_{t}$, we know that $\mathbb{E}_{X_{t}}[Pr\{\hat{X}_{t}\neq X_{t}|X_{t};C_{t}\}]=\mathbb{P}_{t}(C_{t})$, where $\hat{X}_{t}$ denotes the prediction of the channel output transmitted by data symbol $X_{t}$. The online optimization goal for this scenario is to utilize the channel output samples $\{\mathbf{y}_{t}^{j}\}_{j\in[M]},\forall t\in[T]$ transmitted by $M$ codewords in the training sequence for selecting the codebook $C_{t}\in\mathcal{C}\subset(\mathbb{R}^{d})^{M}$ with minimal expected error probability at each round.
Based on channel output samples transmitted by $M$ codewords in the training sequence related to $Z_{t}$, the symbol error rate of codebook $C_{t}$ can be defined as
$\ell_{t}(C_{t}):=\frac{1}{M}\sum_{j=1}^{M}\mathbbm{1}\Big{\{}\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}^{2}<\|\mathbf{x}^{j}_{t}-\mathbf{y}_{t}^{j}\|_{2}^{2}\Big{\}}.$ (9)
Similarly, the transceiver can make use of this symbol error rate as the loss function to perform the online optimization procedure for learning the codebook $C_{t}$ to enhance the transmission performance of data symbols. We devise the multi-armed bandit algorithm to learn $C_{t}$ in Section V.